how many liters should be drained and replaced with pure antifreeze from a radiator containing 12 liters of 40 percent antifreeze to increase the concentration to 50 percent antifreeze.
Answer by stanbon(75887)

how many liters should be drained and replaced with pure antifreeze from a radiator containing 12 liters of 40 percent antifreeze to increase the concentration to 50 percent antifreeze.
----
alcohol - alcohol + alcohol = alcohol
0.40*12 - 0.40x + 1.0x = 0.50*12
-----
Multiply thru by 100 to get:
40*12 - 40x + 100x = 50*12
60x = 10*12
x = 2 liters (amt. to be replaced by pure antifreeze)
==============
Cheers,
Stan H.